# Extragradient methods with complexity guarantees for hierarchical variational inequalities

### v1

@@ 5 Extensions
This section contains several extensions. First, we develop estimates for the case in which the upper level $\operatorname{HVI}$ is strongly monotone, implying that the entire problem (P) has a unique solution. We then move on by explaining how Algorithm 1 can be extended to the Bregman setting, and Tseng’s modified extragradient method.
5.1 Improved rates under stronger monotonicity assumptions
On top the already formulated hypothesis, we impose the following:
@@ 5.1 Improved rates under stronger monotonicity assumptions
Since we assume that the over hierarchical equilibrium problem has a nonempty feasible set, Assumption 4 implies that $\mathcal{S}_{1}=\{z^{*}\}$ for some $z^{*}\in\mathcal{S}_{2}$. We again establish rates on gap functions in terms of the suitably adapted ergodic average
$\bar{z}^{K}=\frac{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}z^{i+1/2}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}},$ (5.1)
where $\gamma_{k}\triangleq\frac{1}{\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)}$ and $\gamma_{0}\triangleq 1$.
@@
Consider problem (P). Let Assumptions 1, 2, 4 be in place. Let additionally either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Let $(\sigma_{k})_{k\geq 0}$ be non-increasing and the sequences $(t_{k})_{k\geq 0}$ and $(\sigma_{k})_{k\geq 0}$ satisfy
$4t^{2}_{k}L^{2}_{k}+2t_{k}\sigma_{k}\mu\leq 1,\qquad k\geq 1.$ (5.2)
@@
$B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.4)
In particular, if the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, then
$B_{\mathcal{U}_{1}}\left[\frac{\sigma_{1}\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}+\frac{\tilde{C}_{\mathcal{U}_{1}\cap\mathcal{S}_{2}}\sum_{i=1}^{k}t_{i}\sigma^{2}_{i}\gamma_{i}}{(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.5)
@@
We start with a refined version of the recursion (4.12). Recall the notation (4.9) and the bound (4.1) that holds for any $z\in\mathcal{Z}$ and $k\geq 1$:
$E_{k+1}(z)$ $\leq E_{k}(z)-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
@@
$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.7)
Using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping (LABEL:eq:nonexpansive), and the step size assumption (5.2), we further obtain
$2t^{2}_{k}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$ $\leq 2t^{2}_{k}t_{k-1}^{2}L_{k}^{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}\leq(1-t_{k}\sigma_{k}\mu)\frac{t_{k-1}^{2}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}.$ (5.8)
Combining (5.7), (5.8), and using the notation (4.11), we obtain
$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq(1-t_{k}\sigma_{k}\mu)D_{k}-D_{k+1}+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.9)
By the strong monotonicity, we obtain
@@
$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\sigma_{k}\mu\norm{z^{k+1/2}-z}^{2}$
$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\frac{\sigma_{k}\mu}{2}\norm{z^{k}-z}^{2}-\sigma_{k}\mu\norm{z^{k+1/2}-z^{k}}^{2},$ (5.10)
where the last inequality follows from the triangle inequality and $(a+b)^{2}/2\leq a^{2}+b^{2}\Leftrightarrow a^{2}\geq(a+b)^{2}/2-b^{2}$. Plugging (5.9) and (5.10) multiplied by 1-1 into (5.6), recalling the notation (4.9) and $G_{k}(\cdot)=g_{2}(\cdot)+\sigma_{k}g_{1}(\cdot)$, we obtain
$E_{k+1}(z)$ $+D_{k+1}\leq(1-t_{k}\sigma_{k}\mu)(E_{k}(z)+D_{k})+(2t^{2}_{k}L^{2}_{k}+t_{k}\mu\sigma_{k}-\frac{1}{2})\norm{z^{k+1/2}-z^{k}}^{2}$
$-t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\right)$
@@
$-t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\right)$ (5.11)
where the last inequality uses (5.2).
We now use this recursion to establish rates on the feasibility and optimality gap, starting with the feasibility gap.
Since the analysis for the monotone setting holds also for strongly monotone case, we have that either Lemma 4.3 still holds under the same assumptions made there (i.e. either Assumption 3 or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ compact). Thus, in any case, for a compact subset $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$, there exists a constant $\tilde{C}_{\mathcal{U}_{2}}$ such that
$\langle{\mathsf{F}}_{1}(z),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq\tilde{C}_{\mathcal{U}_{2}},\qquad\forall z\in\mathcal{U}_{2},k\geq 1.$
Combining this with (5.11), we get
$t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
Multiplying both sides by $\sigma_{k}\gamma_{k}$ (see (5.1) for the definition of $\gamma_{k}$) and using that $\sigma_{k}\geq\sigma_{k+1}$, we obtain
$t_{k}\sigma_{k}\gamma_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq\sigma_{k}\gamma_{k-1}(E_{k}(z)+D_{k})-\sigma_{k+1}\gamma_{k}(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}^{2}\gamma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
Thus, telescoping these inequalities and using the convexity of $g_{2}$, for the ergodic average (5.1), we obtain
$\langle{\mathsf{F}}_{2}(z),\bar{z}^{K}-z\rangle+g_{2}(\bar{z}^{K})-g_{2}(z)$ $\leq\frac{\sigma_{1}(E_{1}(z)+D_{1})+\widetilde{C}_{\mathcal{U}_{2}}\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
Taking the supremum over $\mathcal{U}_{2}$ on both sides and using the definition of the localized gap function (2.4) specialized as the feasibility gap (2.6) together with $D_{1}=0$, we obtain the upper bound (5.3).
@@
$t_{k}\sigma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1}).$
@@
$t_{k}\sigma_{k}\gamma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$ $\leq\gamma_{k-1}(E_{k}(z)+D_{k})-\gamma_{k}(E_{k+1}(z)+D_{k+1}).$
@@
$\langle{\mathsf{F}}_{1}(z),\bar{z}^{K}-z\rangle+g_{1}(\bar{z}^{K})-g_{1}(z)$ $\leq\frac{E_{1}(z)}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
@@
If additionally the lower level solution set enjoys $(\alpha,\rho)$ weak-sharpness we obtain (5.5) by combining (5.4), (2.9), and (5.3). $\blacksquare$
@@
$-B_{\mathcal{U}_{1}}\left[\frac{8L_{{\mathsf{F}}_{2}}(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{\mu K(\alpha/\rho)}+\tilde{C}_{\mathcal{U}_{2}}\frac{4L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu K(\alpha/\rho)}\right]^{1/\rho}\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})$ (5.14)
$\leq\frac{2(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{K},$ (5.15)
where the inequality ()(*) holds under the $(\alpha,\rho)$-weakly sharpness assumption.
@@
First, we show that (5.2) is indeed satisfied:
$4t_{k}^{2}L_{k}^{2}+2t_{k}\sigma_{k}\mu=4\frac{L_{k}^{2}}{16(L_{k}+\sigma_{k}\mu)^{2}}+\frac{2\sigma_{k}\mu}{4(L_{k}+\sigma_{k}\mu)}\leq\frac{1}{4}+\frac{1}{2}\leq 1.$ (5.16)
Further,
$t_{i}\sigma_{i}=\frac{1}{4(L_{{\mathsf{F}}_{2}}+\sigma_{i}(L_{{\mathsf{F}}_{1}}+\mu))}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{L_{{\mathsf{F}}_{2}}+\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{i+\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{1}{\mu}=\frac{1}{i+\varkappa}\cdot\frac{1}{\mu},$ (5.17)
where we denoted $\varkappa=\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)$. This gives us
$\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\prod_{i=1}^{k}(1-\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\mu)=\prod_{i=1}^{k}(1-\frac{1}{i+\varkappa})=\prod_{i=1}^{k}(\frac{i+\varkappa-1}{i+\varkappa})=\frac{\varkappa}{k+\varkappa}$ (5.18)
and $\gamma_{k}=1/\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\frac{k+\varkappa}{\varkappa}$. Finally, we have
$\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}=\sum_{i=1}^{K}\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\frac{i+\varkappa}{\varkappa}=\frac{K}{\mu\varkappa}=\frac{K}{4(L_{{\mathsf{F}}_{1}}+\mu)},$ (5.19)
$\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}=\sum_{i=1}^{K}\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}\cdot\frac{i+\varkappa}{\varkappa}=\frac{L_{{\mathsf{F}}_{2}}}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}\sum_{i=1}^{K}\frac{1}{i}\leq\frac{L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}.$ (5.20)

### v2

@@ 5 Extensions
In this section, first we develop estimates for the case in which the upper-level $\operatorname{HVI}$ is strongly monotone, implying that the entire problem (P) has a unique solution. Second, we show how our method can be extended in a straightforward manner to the forward-backward-forward splitting method of [62].
5.1 Improved rates under strong monotonicity assumption
On top of the already formulated hypothesis, we impose the following:
@@ 5.1 Improved rates under strong monotonicity assumption
Since we assume that the whole hierarchical equilibrium problem (P) has a nonempty feasible set, Assumption 6 implies that $\mathcal{S}_{1}=\{z^{*}\}$ for some $z^{*}\in\mathcal{S}_{2}$. We again establish rates on gap functions in terms of the suitably adapted ergodic average
$\bar{z}^{K}=\frac{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}z^{i+1/2}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}},\text{ where }\gamma_{k}\triangleq\frac{1}{\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)},\gamma_{0}\triangleq 1.$ (5.1)
@@
Consider problem (P). Let Assumptions 1, 2, 4, 6 hold. Let additionally either Assumption 3 hold, or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Let $(\bar{z}^{k})_{k\geq 1}$ as in (5.1) be generated by Algorithm 1 with
$8t^{2}_{k}L^{2}_{k}+2t_{k}\sigma_{k}\mu\leq 1,\qquad k\geq 1.$ (5.2)
@@
$B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{S}_{2})\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.4)
If additionally the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp and $\mathcal{U}_{2}$ is s.t. $\mathcal{U}_{2}\cap\mathcal{S}_{2}\neq\varnothing$, then for the same set $\mathcal{U}_{1}$
$\begin{split}-B_{\mathcal{U}_{1}}&\left[\frac{\sigma_{1}\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}+\frac{\tilde{C}_{\mathcal{U}_{2}}\sum_{i=1}^{k}t_{i}\sigma^{2}_{i}\gamma_{i}}{(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}\right]^{\frac{1}{\rho}}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2}).\end{split}$ (5.5)
@@
We start with a refined version of the recursion (4.5). Recall the bound (4.4) that implies for any $z\in\mathcal{Z}$ and $k\geq 1$:
$E_{k+1}(z)$ $\leq E_{k}(z)-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+D_{k+1}$
@@
$D_{k+1}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1}$
$+\left(4t_{k}^{2}L^{2}_{k}-\frac{1}{2}\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.7)
Acting in the same way as for the proof of (4.13), but with the new step size assumption (5.2) that implies $4t_{k}^{2}L_{k}^{2}\leq 1-t_{k}\sigma_{k}\mu$, we further obtain
$4t^{2}_{k}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$ $\leq 4t^{2}_{k}L_{k}^{2}t_{k-1}^{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}$
$\leq(1-t_{k}\sigma_{k}\mu)D_{k}.$ (5.8)
Combining (5.7), (5.8), we obtain
$D_{k+1}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq(1-t_{k}\sigma_{k}\mu)D_{k}-D_{k+1}+\left(4t_{k}^{2}L^{2}_{k}-\frac{1}{2}\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.9)
By the strong monotonicity of ${\mathsf{F}}_{1}$, we obtain
@@
$\hskip-42.50061pt\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\sigma_{k}\mu\norm{z^{k+1/2}-z}^{2}$
$\hskip-42.50061pt\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle$
$\hskip-34.00049pt+\frac{\sigma_{k}\mu}{2}\norm{z^{k}-z}^{2}-\sigma_{k}\mu\norm{z^{k+1/2}-z^{k}}^{2},$ (5.10)
where the last inequality follows from the triangle inequality and $(a+b)^{2}/2\leq a^{2}+b^{2}\Leftrightarrow a^{2}\geq(a+b)^{2}/2-b^{2}$. Plugging (5.9) and (5.10) multiplied by 1-1 into (5.6), recalling $G_{k}=g_{2}+\sigma_{k}g_{1}$, we obtain
$E_{k+1}(z)$ $+D_{k+1}$
$\leq(1-t_{k}\sigma_{k}\mu)(E_{k}(z)+D_{k})+(4t^{2}_{k}L^{2}_{k}+t_{k}\mu\sigma_{k}-\frac{1}{2})\norm{z^{k+1/2}-z^{k}}^{2}$
$-t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$
$-t_{k}\sigma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$
@@
$-t_{k}\left(H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)+\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)\right),$ (5.11)
where the last inequality uses (5.2) and (2.5). We now use this recursion to establish rates on the feasibility and optimality gap, starting with the feasibility gap.
Since the analysis for the monotone setting holds also for strongly monotone case, we have that either Assumption 3 holds and implies by Lemma 4.2 that the sequences generated by the algorithm are bounded or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact. Thus, in any case, for a compact subset $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$, there exists a constant $\tilde{C}_{\mathcal{U}_{2}}$ such that, for all $z\in\mathcal{U}_{2},k\geq 1$, (cf. the derivation of (4.17))
$H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)=\langle{\mathsf{F}}_{1}(z),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq\tilde{C}_{\mathcal{U}_{2}}.$
Combining this with (5.11) and recalling the notation (4.14), we get
$t_{k}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)W_{k}(z)-W_{k+1}(z)+t_{k}\sigma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
Multiplying both sides by $\sigma_{k}\gamma_{k}$ (see (5.1) for the definition of $\gamma_{k}$) and using that $\sigma_{k}\geq\sigma_{k+1}$ and $W_{k}(z)\geq 0$ for $k\geq 1$, we obtain
$t_{k}\sigma_{k}\gamma_{k}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)$ $\leq\sigma_{k}\gamma_{k-1}W_{k}(z)-\sigma_{k+1}\gamma_{k}W_{k+1}(z)+t_{k}\sigma_{k}^{2}\gamma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
Telescoping these inequalities and using the convexity of $g_{2}$ and $D_{1}=0$, $\gamma_{0}=1$, for the ergodic average (5.1), we obtain
$H^{({\mathsf{F}}_{2},g_{2})}(\bar{z}^{K},z)$ $\leq\frac{\sigma_{1}E_{1}(z)+\widetilde{C}_{\mathcal{U}_{2}}\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
Taking the supremum over $z\in\mathcal{U}_{2}$ on both sides and using the definition of the feasibility gap (2.7), we obtain the upper bound (5.3).
@@
$t_{k}\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1}).$
@@
$t_{k}\sigma_{k}\gamma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)$ $\leq\gamma_{k-1}(E_{k}(z)+D_{k})-\gamma_{k}(E_{k+1}(z)+D_{k+1}).$
@@
$H^{({\mathsf{F}}_{1},g_{1})}(\bar{z}^{K},z)$ $\leq\frac{E_{1}(z)}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
@@
If additionally the lower level solution set enjoys $(\alpha,\rho)$-weak sharpness we obtain (5.5) by combining (5.4), (2.10), and (5.3). $\blacksquare$
@@
$-B_{\mathcal{U}_{1}}$ $\left[\frac{8L_{{\mathsf{F}}_{2}}(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{\mu K(\alpha/\rho)}+\tilde{C}_{\mathcal{U}_{2}}\frac{4L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu K(\alpha/\rho)}\right]^{1/\rho}$ (5.14)
$\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{2(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{K},$
where the inequality ()(*) holds under the $(\alpha,\rho)$-weak sharpness assumption.
@@
First, it is easy to check that (5.2) is indeed satisfied. Further, $t_{i}\sigma_{i}=\frac{1}{L_{{\mathsf{F}}_{2}}+\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{i+\kappa}\cdot\frac{1}{\mu}$, where we denoted $\kappa=\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)$. This gives us
$\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\prod_{i=1}^{k}\left(\frac{i+\kappa-1}{i+\kappa}\right)=\frac{\kappa}{k+\kappa}$
and thus $\gamma_{k}=1/\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\frac{k+\kappa}{\kappa}$. Finally, we have
$\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}$ $=\sum_{i=1}^{K}\frac{1}{i+\kappa}\cdot\frac{1}{\mu}\cdot\frac{i+\kappa}{\kappa}=\frac{K}{\mu\kappa}=\frac{K}{4(L_{{\mathsf{F}}_{1}}+\mu)},$
$\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}$ $=\sum_{i=1}^{K}\frac{1}{i+\kappa}\cdot\frac{1}{\mu}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}\cdot\frac{i+\kappa}{\kappa}=\frac{L_{{\mathsf{F}}_{2}}\sum_{i=1}^{K}\frac{1}{i}}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}\leq\frac{L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}.$